OpenStudy (anonymous):

In early 2012, the price of gas was $3.64. This represents a 40% increase from early 2009. What was the price of gas in early 2009?

OpenStudy (anonymous):

40% of 3.64 = 1.456 Increase from 2009 to2012 , therefore 2009 was cheaper that 2012 therefore: 3.64-1.456 = 2.184

OpenStudy (anonymous):

because if 2012 was an increase from 2009, then 2009 was cheaper.

OpenStudy (anonymous):

Groupon is offering a weight-loss program for $58.50. They says this represents a 55% discount. What was the original value of the weight-loss program?

OpenStudy (anonymous):

55% of 58.50 Get that % answer and then ADD it to the 58.50

OpenStudy (anonymous):

55% of 58.50 = 32.175? So 32.175+58.50=90.675 or 90.68?

OpenStudy (anonymous):

Latex paint sells for $23 per gallon. It will cost you $253 to paint your house. If each gallon will cover 250 square feet, how many square feet is your house?

OpenStudy (anonymous):

The American flag has a width of 15 feet and a length of 5 feet less than twice the width. Determine the amount of square feet of fabric required to make an American flag?

OpenStudy (anonymous):

253 / 23 = 11 gallons. 11 * 250 = 2750sq feet 15 * 2 = 30 - 5 = 25 feet 25 * 15 = 375sq feet

OpenStudy (dwelsher):

x+.4x=3.64 1.4x=3.64 x=3.64/1.4 x=2.60